# A Scalable Measure of Loss Landscape Curvature for Analyzing the Training Dynamics of LLMs

### v1

@@ 1 Introduction
Understanding the evolution of the loss landscape $L(\bm{\theta})$ over the high-dimensional parameter space $\bm{\theta}$ is fundamental to analyzing the training dynamics of neural networks. The loss landscape describes how the objective function changes with model parameters, and its geometry directly influences optimization, generalization and stability throughout training (Gilmer et al., 2022; Kalra and Barkeshli, 2024). Intuitively, gradient-based optimization methods can swiftly navigate to the minima in smooth landscapes, while for rough landscapes, such methods can get trapped, hindering convergence or potentially converging to suboptimal solutions (Li et al., 2018).
@@
The local geometry of the loss landscape is commonly examined through the eigenvalues and eigenvectors of its Hessian matrix $H(\bm{\theta}):=\nabla_{\theta}^{2}L(\bm{\theta})$ (Wu et al., 2018; Lewkowycz et al., 2020; Cohen et al., 2021). These eigenvalues provide insights into the local curvature of the loss landscape —large eigenvalues correspond to sharp, steep directions, while small eigenvalues indicate flat, smooth regions. The largest eigenvalue of the Hessian, which is often referred to as Hessian sharpness $\lambda^{H}_{\max}$, quantifies the worst-case curvature of the landscape and is a fundamental metric in optimization. Its reciprocal, flatness $1/\lambda_{\max}^{H}$, is a complementary measure used to describe the curvature. In neural network optimization, Hessian sharpness relates to the local training stability (Wu et al., 2018; Lewkowycz et al., 2020). For vanilla Gradient Descent (GD), if the learning rate exceeds the threshold $\sim 2/{\lambda_{\max}^{H}}$, loss increases as training ‘catapults’ out of the local basin and eventually converges to a flatter region where the stability conditions are satisfied (Lewkowycz et al., 2020). This observation generalizes to more complex optimization problems, including mini-batch settings and adaptive optimizers, albeit governed by different notions of sharpness (Wu et al., 2018; Agarwala and Pennington, 2025; Cohen et al., 2024; Kalra and Barkeshli, 2024).
@@
Figure 1: Hessian sharpness exhibits progressive sharpening and Edge of Stability (EoS) under constant learning rate. The dashed lines corresponding to the learning rate mark the EoS threshold.
@@
As illustrated in Figure˜1, Hessian sharpness exhibits several robust trends throughout neural network training, particularly with constant learning rates. Training typically begins with an early reduction in Hessian sharpness (Kalra and Barkeshli, 2023; Kalra et al., 2025) followed by a continuous increase until it reaches the stability threshold (Jastrzebski et al., 2020; Cohen et al., 2021). Once this threshold is reached, the training stabilizes through a self-stabilization mechanism (Damian et al., 2023; Cohen et al., 2025) —instead of diverging as classical optimization would predict, Hessian sharpness begins to oscillate around this critical value. The continual increase in Hessian sharpness is referred to as progressive sharpening while the subsequent oscillations around the critical threshold is termed the Edge of Stability (EoS) (Cohen et al., 2021). For the more realistic setting of learning rate schedules involving warmup and decay, sharpness closely follows the learning rate schedule (Gilmer et al., 2022; Kalra and Barkeshli, 2023; Cohen et al., 2021). Due to its particularly close relationship with learning rate, sharpness can be used as a diagnostic tool for identifying training instabilities.
@@
Beyond training dynamics, sharpness of the final solutions is linked to generalization properties, motivated by the intuition that flatter minima generalize better (Hochreiter and Schmidhuber, 1997). However, this relationship has been called into question (Kaur et al., 2023), as empirical analyses show mixed results —while flatter solutions generalize better in some settings, they can hurt performance in others. Sharpness also influences the early phase of training, with flatter initializations typically achieving better performance (Dauphin and Schoenholz, 2019; Kalra and Barkeshli, 2024).
@@
Despite its value for examining model initialization, training dynamics, and generalization properties, analyzing sharpness at scale is challenging. Computing Hessian sharpness relies on iterative eigenvalue solvers (e.g., Power iteration, Lanczos), which require repeated Hessian-vector products (HVPs). However, HVP computation via automatic differentiation is often incompatible with modern training efficiency tools; kernels like Flash Attention (Dao et al., 2022) typically lack second-derivative implementations required for double backpropagation. Furthermore, iterative solvers can require hundreds of iterations with potential convergence failures. As a result, most existing studies are restricted to small-scale experiments (typically $\sim 10$M parameters), leaving open questions about how sharpness evolves in Large Language Models (LLMs) at scale, and how it relates to optimization and downstream performance. In this work, we address this challenge by analyzing critical sharpness, which leverages the relationship between curvature and training instability to serve as a computationally efficient proxy for the loss landscape curvature of LLMs.
@@ 1.1 Contributions
  • We demonstrate that critical sharpness reliably captures well-documented Hessian sharpness phenomena, such as progressive sharpening and the Edge of Stability (Cohen et al., 2021). This makes critical sharpness a practical and computationally efficient proxy for sharpness. Leveraging OLMo-2 checkpoints (Walsh et al., 2025) across pre-training and mid-training, we demonstrate progressive sharpening persists at realistic scales throughout training, including models with up to 7 billion parameters.
  • We introduce a relative measure of critical sharpness $\lambda_{c}^{1\to 2}$ to quantify the curvature of one loss relative to another, with application to analyzing the pre-training loss landscape during mid-training. By varying the mix ratio of pre-training and fine-tuning (e.g. math), we show that increasing the proportion of pre-training data lowers relative critical sharpness, helping the model stay near the “pre-trained basin” during finetuning. We further demonstrate that downstream performance depends on basin retention: GSM8K (Cobbe et al., 2021) (math) benefits from leaving the pre-trained basin ($\eta>2/\lambda_{c}^{1\to 2}$), while MMLU (Hendrycks et al., 2021) (generic reasoning) performs better when the basin is retained ($\eta<2/\lambda_{c}^{1\to 2}$). This enables us to prescribe data mixing strategies to balance performance on math and generic reasoning tasks, depending on the desired outcome.
@@ 2 Characterizing Critical Sharpness: Theory and Empirics
In this section, we characterize the dynamics of critical sharpness both theoretically and empirically. We first establish the theoretical relationship between critical sharpness and Hessian sharpness under a quadratic loss approximation, then validate these insights empirically on FCNs trained on CIFAR-10 using SGD across different batch sizes (Figure˜3).
@@ 2.1 Setup
Figure 2: Comparison of different sharpness measures on an illustrative landscape featuring a sharp valley direction and a flat river direction (Wen et al., 2025). Hessian sharpness quantifies the curvature along the sharpest direction of the landscape (the valley). Directional sharpness measures the quadratic curvature along $\Delta\bm{\theta}$, while critical sharpness, its empirical counterpart, quantifies how far one can step along $\Delta\bm{\theta}$ before the loss increases.
@@ 2.2 Critical Sharpness: A Scalable Measure of Curvature
$\eta_{c}=\min_{\eta>0}\{\eta\mid L(\bm{\theta}-\eta\Delta\bm{\theta})>L(\bm{\theta})\}.$ (1)
@@
To efficiently measure the critical learning rate, we build upon the line search method proposed by Kalra and Barkeshli (2024). Given the update direction $\Delta\bm{\theta}$ from training, we compute the critical learning rate $\eta_{c}$ using a two-phase line search procedure:
@@ 2.3 The Relationship between Critical Sharpness and Hessian Sharpness
Within this approximation, the loss will increase if the learning rate $\eta$ exceeds $2/\lambda_{\text{dir}}$, where $\lambda_{\text{dir}}$ is the directional sharpness (Pan and Li, 2022; Roulet et al., 2024):
Definition 2.2 (Directional Sharpness $\lambda_{\text{dir}}$).
Directional sharpness $\lambda_{\text{dir}}$ serves as an analytically tractable approximation to the empirically measured critical sharpness. Figure˜3 shows that directional sharpness closely tracks critical sharpness throughout training. This alignment suggests that the local quadratic approximation well captures the complex dynamics in this simplistic setting. For Gradient Descent (GD), the directional sharpness can be expressed as a weighted sum of Hessian eigenvalues (Roulet et al., 2024):
@@
For adaptive optimizers (e.g., Adam) with pre-conditioner $P$, the directional sharpness $\lambda_{\text{dir}}$ can be expressed as a weighted sum of the pre-conditioned Hessian eigenvalues $\{\lambda^{PH}_{i}\}_{i=1}^{n}$, where the weights quantify the alignment of the pre-conditioned gradient $P^{-1/2}\bm{g}$ with pre-conditioned Hessian eigendirections $\{\bm{v}_{i}\}_{i=1}^{n}$:
@@
where $c_{i}=P^{-1/2}\bm{g}^{T}\bm{v}_{i}$ is the projection of the pre-conditioned gradient $P^{-1/2}g(\bm{\theta})$ onto the $i^{th}$ eigenvector $\bm{v}_{i}$ of the pre-conditioned Hessian.
@@ 3 Critical Sharpness Dynamics at Scale
Modern large-scale models are typically trained using Adam with weight decay (Loshchilov and Hutter, 2019), which helps mitigate training instabilities (D’Angelo et al., 2024). To analyze sharpness dynamics in this context, we first analyze the stability threshold for common optimizers with weight decay:
@@
We now analyze GPT-style Transformers pre-trained for next-token prediction on the FineWebEdu dataset (Penedo et al., 2024) using AdamW with Warmup-Stable-Decay (WSD) schedule (Hu et al., 2024). Figure˜4 compares the dynamics of critical and pre-conditioned sharpness for three different learning rates. The pre-conditioned sharpness $\lambda^{PH}_{\max}$ exhibits progressive sharpening while following the learning rate schedule throughout training —it is pushed down during the warmup phase, stays constant at the stability threshold during the stable phase, and increases again when the learning rate is decayed. Critical sharpness and directional sharpness follow similar trends throughout training —they exhibit clear progressive sharpening and EoS behavior, while oscillating below the EoS threshold during the stable phase and capturing the increase in pre-conditioned sharpness during the decay phase.
Figure 4: Dynamics of pre-conditioned, directional, and critical sharpness during GPT-style Transformer training on FineWebEdu with AdamW using a Warmup-Stable-Decay (WSD) schedule. Critical sharpness tracks pre-conditioned sharpness throughout training, making it an effective proxy. The colored dashed lines denote the theoretical learning rate threshold $(2+2\beta_{1})/(1-\beta_{1})\eta$ and the two black vertical lines mark the end of warmup and stable phases.
While progressive sharpening has been consistently documented at small scales, especially in full-batch settings, its manifestation and relevance at large scales, particularly in the context of online LLM training, remain largely unexplored. Having established that critical sharpness serves as an efficient and reliable proxy for progressive sharpening and Edge of Stability phenomena, we are now in the position to investigate whether progressive sharpening persists in large-scale LLM training. To this end, we analyze the publicly available OLMo-2 77B checkpoints (Walsh et al., 2025). OLMo-2 provides checkpoints throughout both pre-training and mid-training stages, enabling us to study sharpness dynamics at scale without the computational cost of training such models from scratch.
The OLMo-2 models are trained using a two-stage curriculum approach. During the pre-training stage ($90\%$ of compute), the model is trained for 44T tokens on the DCLM dataset (Li et al., 2025), a diverse web-sourced corpus. Following pre-training, the model undergoes a mid-training stage, where it is trained for 5050B tokens on a curated mix. This mix includes academic papers, Wikipedia articles, instruction-following data, StackExchange documents, and approximately $50\%$ of the original pre-training dataset (DCLM). During the pre-training stage, the model is trained using a learning rate schedule consisting of 2,0002{,}000 steps of warmup, followed by a cosine decay down to one-tenth of its peak value [^2] For the OLMo-2 77B models, the cosine decay schedule is designed to reach one-tenth of the peak learning rate at 55T tokens, but is truncated at 44T tokens.. By comparison, during the mid-training stage, the learning rate is linearly decayed to zero. As the learning rate is continuously decreased throughout training after the warmup stage, the EoS threshold increases, and consequently we expect sharpness to progressively increase for the remainder of training. To confirm this, we examine how the critical sharpness evolves across pre-training and mid-training using the publicly available checkpoints (see Section˜8 for the details). Figure˜5 shows that critical sharpness rapidly decreases during early training, but then continually increases (progressive sharpening) throughout both the pre-training and mid-training stages. This result provides the first empirical evidence of progressive sharpening at scale in practical LLM training settings.

### v2

@@ 1 Introduction
Understanding the evolution of the loss landscape $L(\bm{\theta})$ over the high-dimensional parameter space $\bm{\theta}$ is fundamental to analyzing the training dynamics of neural networks. The loss landscape describes how the objective function changes with model parameters, and its geometry directly influences optimization, generalization, and stability throughout training (gilmer2022a; kalra2024why). Intuitively, gradient-based optimization methods can swiftly navigate to the minima in smooth landscapes, while for rough landscapes, such methods can get trapped, hindering convergence or potentially converging to suboptimal solutions (losslandscapes).
@@
The local geometry of the loss landscape is commonly examined through the eigenvalues and eigenvectors of its Hessian matrix $H(\bm{\theta}):=\nabla_{\theta}^{2}L(\bm{\theta})$ (Wu2018_how_sgd; lewkowycz2020largelearningratephase; cohen2021gradient). These eigenvalues provide insights into the local curvature of the loss landscape —large eigenvalues correspond to sharp, steep directions, while small eigenvalues indicate flat, smooth regions. The largest eigenvalue of the Hessian, which is often referred to as Hessian sharpness $\lambda^{H}_{\max}$, quantifies the worst-case curvature of the landscape and is a fundamental metric in optimization. Its reciprocal, flatness $1/\lambda_{\max}^{H}$, is a complementary measure used to describe the curvature. In neural network optimization, Hessian sharpness relates to the local training stability (Wu2018_how_sgd; lewkowycz2020largelearningratephase). For vanilla Gradient Descent (GD), if the learning rate exceeds the threshold $\sim 2/{\lambda_{\max}^{H}}$, loss increases as training ‘catapults’ out of the local basin and eventually converges to a flatter region where the stability conditions are satisfied (lewkowycz2020largelearningratephase). This observation generalizes to more complex optimization problems, including mini-batch settings and adaptive optimizers, albeit governed by different notions of sharpness (Wu2018_how_sgd; agarwala2025highdimensionalanalysisreveals; cohen2024adaptivegradientmethodsedge; kalra2024why). For adaptive optimizers (e.g., Adam), the relevant measure is the pre-conditioned sharpness $\lambda^{PH}_{\max}$ (cohen2024adaptivegradientmethodsedge), which we formalize in Section˜2.1.
@@
Figure 1: Pre-conditioned Hessian sharpness ($\lambda^{PH}_{\max}$) exhibits progressive sharpening and Edge of Stability (EoS) under constant learning rate. The dashed lines corresponding to the learning rate mark the EoS threshold.
@@
As illustrated in Figure˜1, Hessian sharpness exhibits several robust trends throughout neural network training, particularly with constant learning rates. Training typically begins with an early reduction in Hessian sharpness (kalra2023phase; kalra2025universal) followed by a continuous increase until it reaches the stability threshold (Jastrzebski2020The; cohen2021gradient). Once this threshold is reached, the training stabilizes through a self-stabilization mechanism (damian2023selfstabilization; cohen2025understanding) —instead of diverging as classical optimization would predict, Hessian sharpness begins to oscillate around this critical value. The continual increase in Hessian sharpness is referred to as progressive sharpening while the subsequent oscillations around the critical threshold are termed the Edge of Stability (EoS) (cohen2021gradient). For the more realistic setting of learning rate schedules involving warmup and decay, sharpness closely follows the learning rate schedule (gilmer2022a; kalra2024why; cohen2021gradient). Due to its particularly close relationship with learning rate, sharpness can be used as a diagnostic tool for identifying training instabilities.
@@
Beyond training dynamics, sharpness of the final solutions is linked to generalization properties, motivated by the intuition that flatter minima generalize better (Hochreiter1997FlatM). However, this relationship has been called into question (pmlr-v187-kaur23a), as empirical analyses show mixed results —while flatter solutions generalize better in some settings, they can hurt performance in others. Sharpness also influences the early phase of training, with flatter initializations typically achieving better performance (NEURIPS2019_876e8108; kalra2024why).
@@
Despite its value for examining model initialization, training dynamics, and generalization properties, analyzing sharpness at scale is challenging. Computing Hessian sharpness relies on iterative eigenvalue solvers (e.g., Power iteration, Lanczos), which require repeated Hessian-vector products (HVPs). However, HVP computation via automatic differentiation is often incompatible with modern training efficiency tools; kernels like Flash Attention (dao2022flashattention) typically lack second-derivative implementations required for double backpropagation. Furthermore, iterative solvers can require hundreds of iterations with potential convergence failures. As a result, most existing studies are restricted to small-scale experiments (typically $\sim 10$M parameters), leaving open questions about how sharpness evolves in Large Language Models (LLMs) at scale, and how it relates to optimization and downstream performance. In this work, we address this challenge by analyzing critical sharpness, which leverages the relationship between curvature and training instability to serve as a computationally efficient proxy for the loss landscape curvature of LLMs.
@@ 1.1 Contributions
  • We demonstrate that critical sharpness reliably captures well-documented Hessian sharpness phenomena, such as progressive sharpening and the Edge of Stability (cohen2021gradient). This makes critical sharpness a practical and computationally efficient proxy for sharpness. Leveraging OLMo-2 checkpoints (walsh2025) across pre-training and mid-training, we demonstrate progressive sharpening persists at realistic scales throughout training, including models with up to 7 billion parameters.
  • We introduce a relative measure of critical sharpness $\lambda_{c}^{1\to 2}$ to quantify the curvature of one loss relative to another, with application to analyzing the pre-training loss landscape during mid-training/fine-tuning. By varying the mix ratio of pre-training and fine-tuning (e.g. math), we show that increasing the proportion of pre-training data lowers relative critical sharpness, helping the model stay near the “pre-trained basin” during fine-tuning. We further demonstrate that downstream performance depends on basin retention: GSM8K (cobbe2021training) (math) benefits from leaving the pre-trained basin ($\eta>2/\lambda_{c}^{1\to 2}$), while MMLU (hendryckstest2021) (generic reasoning) performs better when the basin is retained ($\eta<2/\lambda_{c}^{1\to 2}$). This enables us to prescribe data mixing strategies to balance performance on math and generic reasoning tasks, depending on the desired outcome.
@@ 2 Characterizing Critical Sharpness: Theory and Empirics
In this section, we characterize the dynamics of critical sharpness both theoretically and empirically. We first establish the theoretical relationship between critical sharpness and Hessian sharpness under a quadratic loss approximation, then validate these insights empirically on MLPs trained on CIFAR-10 using SGD across different batch sizes (Figure˜3).
@@ 2.2 Critical Sharpness: A Scalable Measure of Curvature
$\eta_{c}=\inf\{\eta>0\mid L(\bm{\theta}-\eta\Delta\bm{\theta})>L(\bm{\theta})\}.$ (1)
@@
To efficiently measure the critical learning rate, we build upon the line search method proposed by kalra2024why. Given the update direction $\Delta\bm{\theta}$ from training, we compute the critical learning rate $\eta_{c}$ using a two-phase line search procedure:
@@
Figure 2: Comparison of different sharpness measures on an illustrative landscape featuring a sharp valley direction and a flat river direction (wen2025understanding). Hessian sharpness quantifies the curvature along the sharpest direction of the landscape (the valley). Directional sharpness measures the quadratic curvature along $\Delta\bm{\theta}$, while critical sharpness, its empirical counterpart, quantifies how far one can step along $\Delta\bm{\theta}$ before the loss increases.
@@ 2.3 The Relationship between Critical Sharpness and Hessian Sharpness
Within this approximation, the loss will increase if the learning rate $\eta$ exceeds $2/\lambda_{\text{dir}}$, where $\lambda_{\text{dir}}$ is the directional sharpness (pan2022toward; roulet2024stepping):
Definition 2.2 (Directional Sharpness $\lambda_{\text{dir}}$).
Directional sharpness $\lambda_{\text{dir}}$ serves as an analytically tractable approximation to the empirically measured critical sharpness. Figure˜3 shows that directional sharpness closely tracks critical sharpness throughout training. This alignment suggests that the local quadratic approximation well captures the complex dynamics in this simplistic setting. For Gradient Descent (GD), the directional sharpness can be expressed as a weighted sum of Hessian eigenvalues (roulet2024stepping):
@@
For adaptive optimizers with pre-conditioner $P$ and update direction $P^{-1}\bm{g}$ (e.g., RMSProp), the directional sharpness $\lambda_{\text{dir}}$ can be expressed as a weighted sum of the pre-conditioned Hessian eigenvalues $\{\lambda^{PH}_{i}\}_{i=1}^{n}$, where the weights quantify the alignment of the pre-conditioned gradient $P^{-1/2}\bm{g}$ with pre-conditioned Hessian eigendirections $\{\bm{v}_{i}\}_{i=1}^{n}$:
@@
where $c_{i}=P^{-1/2}\bm{g}^{T}\bm{v}_{i}$ is the projection of the pre-conditioned gradient $P^{-1/2}g$ onto the $i^{th}$ eigenvector $\bm{v}_{i}$ of the pre-conditioned Hessian. For momentum-based optimizers like Adam, this result serves as an approximation.
@@ 3 Critical Sharpness Dynamics at Scale
Figure 4: Dynamics of pre-conditioned, directional, and critical sharpness during GPT-style Transformer training on FineWebEdu with AdamW using a Warmup-Stable-Decay (WSD) schedule. Critical sharpness tracks pre-conditioned sharpness throughout training, making it an effective proxy. The colored dashed lines denote the theoretical learning rate threshold $\left(\frac{2}{\eta}-\gamma\right)\left(\frac{1+\beta_{1}}{1-\beta_{1}}\right)$ and the two black vertical lines mark the end of warmup and stable phases.
Modern large-scale models are typically trained using Adam with weight decay (adamwloshchilov2018), which helps mitigate training instabilities (dangelo2024why). To analyze sharpness dynamics in this context, we first analyze the stability threshold for common optimizers with weight decay:
@@
We now analyze GPT-style Transformers pre-trained for next-token prediction on the FineWebEdu dataset (penedo2024the) using AdamW with Warmup-Stable-Decay (WSD) schedule (hu2024minicpm). Figure˜4 compares the dynamics of critical and pre-conditioned sharpness for three different learning rates. The pre-conditioned sharpness $\lambda^{PH}_{\max}$ exhibits progressive sharpening while following the learning rate schedule throughout training —it is pushed down during the warmup phase, stays constant at the stability threshold during the stable phase, and increases again when the learning rate is decayed. Critical sharpness and directional sharpness follow similar trends throughout training —they exhibit clear progressive sharpening and EoS behavior, while oscillating below the EoS threshold during the stable phase and capturing the increase in pre-conditioned sharpness during the decay phase.
While progressive sharpening has been consistently documented at small scales, especially in full-batch settings, its manifestation and relevance at large scales, particularly in the context of online LLM training, remain largely unexplored. Having established that critical sharpness serves as an efficient and reliable proxy for progressive sharpening and Edge of Stability phenomena, we are now in the position to investigate whether progressive sharpening persists in large-scale LLM training. To this end, we analyze the publicly available OLMo-2 77B checkpoints (walsh2025). OLMo-2 provides checkpoints throughout both pre-training and mid-training stages, enabling us to study sharpness dynamics at scale without the computational cost of training such models from scratch.
@@
The OLMo-2 models are trained using a two-stage curriculum approach. During the pre-training stage ($90\%$ of compute), the model is trained for 44T tokens on the DCLM dataset (li2025datacomplmsearchgenerationtraining), a diverse web-sourced corpus. Following pre-training, the model undergoes a mid-training stage, where it is trained for 5050B tokens on a curated mix. This mix includes academic papers, Wikipedia articles, instruction-following data, StackExchange documents, and approximately $50\%$ of the original pre-training dataset (DCLM). During the pre-training stage, the model is trained using a learning rate schedule consisting of 2,0002{,}000 steps of warmup, followed by a cosine decay down to one-tenth of its peak value [^2] For the OLMo-2 77B models, the cosine decay schedule is designed to reach one-tenth of the peak learning rate at 55T tokens, but is truncated at 44T tokens.. By comparison, during the mid-training stage, the learning rate is linearly decayed to zero. As the learning rate is continuously decreased throughout training after the warmup stage, the EoS threshold increases, and consequently we expect sharpness to progressively increase for the remainder of training. To confirm this, we examine how the critical sharpness evolves across pre-training and mid-training using the publicly available checkpoints (see Section˜8 for the details). Figure˜5 shows that critical sharpness rapidly decreases during early training, but then continually increases (progressive sharpening) throughout both the pre-training and mid-training stages. This result provides the first empirical evidence of progressive sharpening at scale in practical LLM training settings.